OpenStudy (anonymous):

The sum of the interior angles of a regular polygon is 1800 degrees. What is the measure of each exterior angle of the polygon? A. 150 degrees B. 108 degrees C. 60 degrees D. 30 degrees

OpenStudy (anonymous):

First you need to find how many sides the polygon has, so you use: Sum of the measures of the interior angles of a polygon formula = 180(n - 2) With n being number of sides Plug in what you know 1800 = 180(n - 2) Solve 1800 = 180n - 360 2160 = 180n 12 = n n = 12 Next you find the measure of one interior angle, using m=  With m being the measure of an interior angle Plug in what you know m = 180(12 - 2)/12 Solve m = 180(10)/12 m = 1800/12 m = 150 So each interior angle = 150 degrees. Since an interior angle + an exterior angle = 180 degrees, just subtract: 180 - 150 ----- 30 So the answer ends up being 30 degrees.